OpenStudy (anonymous):

Help please! >< how do i prove that if the polynomial f(x) has a factor (x-a), then f(a)=0.

OpenStudy (lalaly):

f(x) = (x - a)(another polynomial say h(x)) f(x) = (x - a)*h(x) f(a) = (a - a)*h(a) f(a) = 0*h(a) f(a) = 0

OpenStudy (anonymous):

thank you!